ann is 4 years older than toby bill is twice toby's age if the sum of their ages is 28 how old is ann ?
Found 2 solutions by josgarithmetic, ikleyn:
Answer by josgarithmetic(39618)

-----------------------------------------------------------------------
ann is 4 years older than toby bill is twice toby's age if the sum of their ages is 28 how old is ann ?
---------------------------------------------------------------------

THIS is what you have:
------------------------------------------------------------
Ann is 4 years older than Toby. Bill is twice toby's age. If the sum of their ages is 28, how old is Ann ?
------------------------------------------------------------

Answer by ikleyn(52794)

.
Ann is 4 years older than Toby.
Bill is twice Toby's age
if the sum of their ages is 28, how old is Ann ?
~~~~~~~~~~~~

Let x be Ann age.


Then the Toby's age is (x-4) years and Bill is  2(x-4) yearts old.


The sum is 28 years


    x + (x-4) + 2(x-4) = 28

    (x-4) + (x-4) + 2(x-4) = 28-4

    4(x-4) = 24

      x-4  = 24/4 = 6

       x          = 6 + 4 = 10.


ANSWER.  Ann is 10 years old.

Solved.